OpenStudy (anonymous):

The time t required to drive a certain distance varies inversely with the speed r. If it takes 4 hours to drive the distance at 40 miles per hour, how long will it take to drive the same distance at 55 miles per hour?

OpenStudy (emmamink):

ok hold on

OpenStudy (anonymous):

ok

OpenStudy (emmamink):

@TheSmartOne

TheSmartOne (thesmartone):

Hi, do you know what inversely proportional means?

OpenStudy (anonymous):

no.

OpenStudy (emmamink):

Omg u r an ambassador too? congrats!

TheSmartOne (thesmartone):

Seem familiar? \(\sf\LARGE y=\frac{k}{x}\)

OpenStudy (anonymous):

no.

OpenStudy (emmamink):

lol im gonna do trig and i dont even recognize that. i am really tired today. cant think straight....

TheSmartOne (thesmartone):

y is inversely proportional to x so then that makes \(\sf\LARGE y=\frac{k}{x}\) where k is a constant

TheSmartOne (thesmartone):

The time t required to drive a certain distance varies inversely with the speed r. That means \(\sf\Large t = \frac{k}{r}\)

TheSmartOne (thesmartone):

If it takes 4 hours to drive the distance at 40 miles per hour. Use that infomation to get the value of k

TheSmartOne (thesmartone):

And then finally, plug back k into that formula and plug in the last piece of information they gave you to solve for time how long will it take to drive the same distance at 55 miles per hour?

OpenStudy (anonymous):

2.91 hours

TheSmartOne (thesmartone):

what's that?

OpenStudy (anonymous):

t = 2.91 hours.

TheSmartOne (thesmartone):

they told you If it takes 4 hours to drive the distance at 40 miles per hour. t there is 4

OpenStudy (anonymous):

oh.

TheSmartOne (thesmartone):

but that isn't your final answer :p

TheSmartOne (thesmartone):

Are you saying that t=2.91 as your final answer? @allen10

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

That is correct :)